We know the following expected returns for stock A and the market portfolio, given different states of the economy:

State (s) Probability E(rA,s) E(rM,s)
Recession 0.3 -0.04 0.04
Normal 0.5 0.11 0.07
Expansion 0.2 0.19 0.11

The risk-free rate is 0.02.

Part 1

Assuming the CAPM holds, what is the beta for stock A?
Answer #1
State (s) Probability E(rA,s) E(rM,s)
Recession 0.3 -0.04 0.04
Normal 0.5 0.11 0.07
Expansion 0.2 0.19 0.11

Expected Return (when probabilities of different states of economies are given) is calculated using the formula:

E[R] = p1*R1 + p2*R2 + p3*R3

Expected return of stock A = E[RA] = 0.3*(-0.04) + 0.5*0.11 + 0.2*0.19 = 8.1%

Expected return on Market = E[RM] = 0.3*0.04 + 0.5*0.07 + 0.2*0.11 = 6.9%

Risk-free rate = 0.02 = 2%

Beta of stock A = βA

According to CAPM, expected return on stock A is give by:

E[RA] = RF + βA*(E[RM] - RF)

8.1% = 2% + βA*(6.9% - 2%)

8.1% - 2% = βA*4.9%

βA = 6.1%/4.9% = 1.2448980

Beta of stock A = 1.2448980